OpenStudy (anonymous):

Solve the system of equations. y=-0.04x^2+8.3x y=1.12x+1.6 a)no solution b) 0.22, 1.85 c) 179.28, 202.39 d) 0.22,1.85 and 179.28, 202.39

OpenStudy (anonymous):

i have no idea how to do this crap. math is not my subject :/

OpenStudy (callisto):

y=-0.04x^2+8.3x -(1) y=1.12x+1.6 (2) (1) - (2) 0 = -0.04x^2+8.3x - (1.12x+1.6) 0 = -0.04x^2 +7.18x - 1.6 Use the quadratic formula \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] Put a = -0.04, b=7.18 and c=-1.6 into the formula, what do you get?

OpenStudy (anonymous):

i donno i dont know how to do it ..... if i know i dtuldnt need the help on here

OpenStudy (callisto):

which part you don't know?

OpenStudy (anonymous):

all of it. and hard to learn when my ebook barely covers it. and fact that its been almost 4 years of doing this stuff

OpenStudy (callisto):

Do you know how to solve simultaneous equations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ive tried and i just dont get it and i here il ive tohee p live to help due to my husbands military and moving and so so no friends here yet

OpenStudy (anonymous):

sorry my keyboards acting up

OpenStudy (callisto):

Never mind. In this case, do you mind starting with learning to solve simple simultaneous equations? It would be better for you to understand it...

OpenStudy (anonymous):

its not going to help me i need like a person to sit with me an work it out blah this is frustrating. brb my sons screaming for me

OpenStudy (anonymous):

back

OpenStudy (callisto):

I don't mind helping you understand on OS though...

OpenStudy (anonymous):

no i mean like person to person , i should of done this ii t hlas but with a son a husband bout to deploy i got stuck doing it online

OpenStudy (anonymous):

in a class*

OpenStudy (callisto):

Hmmm...... How can I help?

OpenStudy (anonymous):

i dont know how u get the answer because thats what my stuff nsupr matches up on is the problem to the answer

OpenStudy (anonymous):

because i got b town i tried to work it out

OpenStudy (anonymous):

answer B

OpenStudy (callisto):

Hold on..

OpenStudy (callisto):

Are you sure the answer is B? I've got different answers from you..

OpenStudy (anonymous):

different answers from me ? thats what i got when i tried it but i donno how to do it is why i asked on here . because i dont know how to do it and its probably not right

OpenStudy (anonymous):

but forget it now cause my darn power went out and wiped it all out so i gotta start my online assignment all over and it changes the problems :/ blah

OpenStudy (callisto):

Sorry, I wasn't here for a while. For this question, there should be 2 pairs of solution. Answer B just showed you one pair of them ..